40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 929     Entry time: Sat Apr 21 00:55:00 2012
Author: frank, tara 
Type: DailyProgress 
Category: BEAT 
Subject: Beat measurement update 

We measured the beat after a few changes in the setup (modulation index, air spring, small enclosure around the beat). We update the noise budget with suppressed laser frequency noise and Marconi's frequency noise in ACAV loop as well. There is not much improvement in the overall beat noise.

A few changes in the setup:

  •  Air Springs: We activated the air springs (it was off because we though it causes the vacuum chamber to move too much during the work), so there is one more seismic isolation stage. The peak at 58 Hz disappears.  However, there are peaks around 6-12 Hz showing up this time. They are probably from the table and the legs that are not set to float the table properly*.
  • Modulation index: now the modulation index is 0.18. I measured the power in the carrier and sideband to calculate the mod index, see PSL:855 . This is the maximum we can have with the 13dBm from Marconi with 4-way splitter (-6dBm),-0.5 dBm loss from the cable, so the power to EOM is 6.5dBm.  With the new mod index, the frequency discriminator gain (inverse) for RCAV/ACAV become 22 and 31 kHz/V.
  • A prototype enclosure box around beat setup: PSL:925.

beat_2012_04_20.png

Updated Traces:

  •   Suppressed Laser frequency noise: Since we changed the OLG of TTFSS by changing the modulation index, we remeasured the OLG of TTFSS, applied to the laser frequency noise (estimated to be 1e4/f Hz/rtHz). Good news, we have enough gain in RCAV loop to suppress the laser noise to coating level. The gain setup for common/fast is 750/750
  •  Frequency noise of Marconi for ACAV: We had not done this so far, so we wanted to make sure that it would not be the limiting source. This is done by dividing the frequency noise of marconi by ACAV OLG TF. We assume that the frequency noise is the same as we measured from the marconi for PLL plotted in the noise budget.
  • Shot noise is updated with the current modulation index (0.18) and cavities visibilities (~80%). It is not shown in the nb because I included it in the PDH detection noise.

* The table legs have small leaks, so we have air compressor to keep the pressure high, PSL:880. It works once every hour when the pressure is low.  The change of pressure can screw up the beam alignment a lot (for example, DC readout from RCAV RFPD can be varied from 100mV - 1.2V, with 1.7V maximum value (off resonance). I think it affects the alignment to AOM and change the diffraction efficiency/ beam shape as well.  It also changes the seismic isolation property. The peaks around 10 Hz show up when the pressure is low.  I'll find new legs from New Port and ask for the quote.

 

==Plan for the next few days==

1) Try using DAQ to measure beat signal, see if we can get rid off the frequency noise from PLL

2) Find out about the new legs(prices/availability)

3) Make the acoustic enclosure box around the whole setup

4) Work on the design for new vacuum windows.

 

ELOG V3.1.3-